OpenStudy (anonymous):

1. What is the area of the circle? Use 3.14 for CE_LEAP_MA_0004_pie symbol. CE_four mm circle (1 point)1. 12.56 mm2 2. 25.12 mm2 3. 50.24 mm2 or 4. 502.4 mm2

OpenStudy (anonymous):

@kirbykirby @mathbrz @mathmale

OpenStudy (anonymous):

Easy

OpenStudy (anonymous):

im just confused ive been getting different answers lately

OpenStudy (anonymous):

Area of circle = \[\pi \times r ^{2}\]

OpenStudy (anonymous):

r = 4

OpenStudy (anonymous):

\[\pi \times (4)^{2}\]

OpenStudy (anonymous):

I just got 502.4

OpenStudy (anonymous):

\[\pi \times 16\]

OpenStudy (anonymous):

\[3.14 \times 16 = 50.24\]

OpenStudy (anonymous):

see what I mean almost the same but mines is wrong

OpenStudy (anonymous):

so, \[Area of Circle = 50.24mm ^{2}\]

OpenStudy (anonymous):

i think you had a calculation error

OpenStudy (anonymous):

with the decimal

OpenStudy (anonymous):

place

OpenStudy (anonymous):

yeah I had a few taking a test just now but I got a 70% on it

OpenStudy (anonymous):

or you used incorrect formulas

OpenStudy (anonymous):

You have to be confident in your ability of solving a problem. But, to be confident you have to do a lot of problems until you have the formula jotted down on your memmory

OpenStudy (anonymous):

true